OpenStudy (anonymous):

the equation y = 1777x - 27,153 can be used to predict the number y of gun deaths in the united states x years after 2000, that is x = 0 corresponds to 2000, x = 3 corresponds to 2003, x = 4 corresponds to 2004, and so on. Predict the number of gun deaths in 2004 and 2009. In what year will the number of gun deaths be 7606? The predicted gun death in 2004 will be --? The predicted gun death in 2009 will be --? The predicted number of gun deaths will be 7606 in the year --?

OpenStudy (anonymous):

The predicted gun death in 2004 will be 1777(4) -27 153 =7108 - 27 153 = - 20 045 which does not make sense. If x = 2004 we get 1777(2004) -27 153 =3 533 955 This is about 1/10 the population af the USA and is off the mark therefore! Have you posted the exact question that was given?

OpenStudy (anonymous):

this is the first why

OpenStudy (anonymous):

Well given the equation and the two approaches I have tried, both give extraordinary answers.

OpenStudy (anonymous):

soeey i messed up on my equation it says the equation y = -1777x + 27,153 can be used to predict the number y of gun deaths in the united states x years after 2000, that is x = 0 corresponds to 2000, x = 3 corresponds to 2003, x = 4 corresponds to 2004, and so on. Predict the number of gun deaths in 2004 and 2009. In what year will the number of gun deaths be 7606? The predicted gun death in 2004 will be --? The predicted gun death in 2009 will be --? The predicted number of gun deaths will be 7606 in the year --?

OpenStudy (anonymous):

Gianfranco sorry i messed up on my equation it says the equation y = -1777x + 27,153 can be used to predict the number y of gun deaths in the united states x years after 2000, that is x = 0 corresponds to 2000, x = 3 corresponds to 2003, x = 4 corresponds to 2004, and so on. Predict the number of gun deaths in 2004 and 2009. In what year will the number of gun deaths be 7606? The predicted gun death in 2004 will be --? The predicted gun death in 2009 will be --? The predicted number of gun deaths will be 7606 in the year --?

OpenStudy (anonymous):

Ok Jazi just look at my first method and it will now work.

OpenStudy (anonymous):

y = -1777x + 27,153 y = -1777(4) + 27,153 = 20 045 Do the same for the other years given Last problem... 7606 = -1777x + 27,153 -19547 = -1777x divide by -1777 on both sides x = 11 therefore the year this will happen is 2011